An airplane travels 1800 miles in 2.5 hours with the aid of a tail wind. It takes 3 hours for the return trip, flying against the same wind. Find the speed of the airplane in the still wind, and the speed of the wind. Airplane speed:___________miles per hour Wind speed:____________miles per hour,
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 3 hours
with wind 2.5 hours

Distance against 1800 miles distance with 1800 miles
t=d/r against wind -
1800 / ( x - y )= 3
3 ( x - y ) = 1800
3 x - 3 y = 1800 ....................1

1800 / ( x + y )= 2.50
2.50 ( x + y ) = 1800
2.50 x + 2.50 y = 1800 ...............2
Multiply (1) by 2 1/2
Multiply (2) by 3
we get
7.50 x + -8 y = 4500
7.50 x + 8 y = 5400
15 x = 9900
/ 15
x = 660 mph

plug value of x in (1)
3 x -3 y = 1800
1980 -3 y = 1800
-3 y = 1800 -1980
-3 y = -180 mph
y = 60
Plane speed 660 mph
wind speed 60 mph

m.ananth@hotmail.ca